# Debiased machine learning for combining probability and non-probability survey data

### v1

@@ 2 Data-generating mechanism and estimator
$\hat{\theta}_{1}(\pi^{B},m)=\frac{1}{n}\sum_{i=1}^{n}U_{i}(\pi^{B},m),$
$U(\pi^{B},m)=U(\pi^{B},m;X,Y,R^{A},R^{B},\pi^{A})=\frac{R^{B}}{\pi^{B}(X)}Y+\left\{\frac{R^{A}}{\pi^{A}}-\frac{R^{B}}{\pi^{B}(X)}\right\}m(X),$
[…]
3 von Mises expansion
3.1 Independent, identically distributed data and infinite population
[…]
$\hat{\theta}_{\rm iid}(\pi_{\rm iid},m_{\rm iid})=\frac{1}{n}\sum_{i=1}^{n}\frac{R_{i}}{\pi_{\rm iid}(X_{i})}Y_{i}+\left\{1-\frac{R_{i}}{\pi_{\rm iid}(X_{i})}\right\}m_{\rm iid}(X_{i}).$
Here, $R$ is an indicator that $Y$ is observed, $\hat{\pi}_{\rm iid}(X)$ is an estimator of $\pi_{\rm iid0}(X)=P(R=1\mid X)$ and $\hat{m}_{\rm iid}(X)$ is an estimator of $m_{\rm iid0}(X)=E(Y\mid X)$. Hines et al. (2022)[20] give an account of the properties of $\hat{\theta}_{\rm iid}$ and of the von Mises expansion that can be used to derive these properties. They describe how, if $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ converge to $\pi_{\rm iid0}$ and $m_{\rm iid0}$ sufficiently rapidly as $n\rightarrow\infty$, then $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})=\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})+o_{p}(n^{-1/2})$, and so $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ has the same asymptotic normal distribution as $\hat{\theta}_{\rm iid}(\pi_{\rm iid0},m_{\rm iid0})$. The required convergence rates of $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are slower than the parametric $n^{1/2}$ rate, which enables the use of data-adaptive estimators. Often, $\hat{\theta}_{\rm iid}$ is used with cross-fitting, a technique similar to cross-validation. This involves splitting the sample of $n$ individuals into subsamples, called folds, calculating estimates $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ using the data on all but fold $k$, and then calculating $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ using these estimates and the data on fold $k$. This procedure is repeated for $k=1,\ldots,K$, and the resulting $K$ values of $\hat{\theta}_{\rm iid}(\hat{\pi}_{\rm iid},\hat{m}_{\rm iid})$ are averaged. This ensures that $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ are independent of the data in fold $k$, which avoids the need for $\hat{\pi}_{\rm iid}$ and $\hat{m}_{\rm iid}$ to satisfy the Donsker condition[20].
3.2 Sampling without replacement from a finite population
In this article, we obtain an analogous asymptotic equivalence result for our sample survey setting. We make use of cross-fitting. Randomly partition the set of $n$ individuals into $K$ (e.g. $K=5$) subsets, called folds. The way in which is done will be described later. Let $\mathcal{S}_{k}\subset\{1,\ldots,n\}$ denotes the set of indices of the individuals belonging to fold $k$ ($k=1,\ldots,K$). Let $n_{k}=|\mathcal{S}_{k}|$ denote the number of individuals in $\mathcal{S}_{k}$. Let $\mathcal{S}_{-k}$ denote the indices of all individuals except those in fold $k$ (so, $\mathcal{S}_{k}\cup\mathcal{S}_{-k}=\{1,\ldots,n\}$). Let $\hat{m}_{k}$ be an estimator of $m_{0}$ calculated using only the data $\{(X_{i},Y_{i}):R^{B}_{i}=1\mbox{ and }i\in\mathcal{S}_{-k}\}$. Let $\hat{\pi}^{B}_{k}$ be an estimator of $\pi^{B}_{0}$ obtained using only the data $\{(R^{A}_{i},R^{B}_{i}):i\in\mathcal{S}_{-k}\}$ and $\{X_{i}:R^{A}_{i}+R^{B}_{i}\geq 1\mbox{ and }i\in\mathcal{S}_{-k}\}$. Write $\underline{\pi}^{B}=(\pi^{B}_{1},\ldots,\pi^{B}_{K})$, $\underline{m}=(m_{1},\ldots,m_{K})$, $\underline{\hat{\pi}}^{B}=(\hat{\pi}^{B}_{1},\ldots,\hat{\pi}^{B}_{K})$ and $\underline{\hat{m}}=(\hat{m}_{1},\ldots,\hat{m}_{K})$. In a slight abuse of notation, we define
$\hat{\theta}_{1}(\underline{\pi}^{B},\underline{m})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\pi^{B}_{k},m_{k}).$
[…]
@@ 3.2 Sampling without replacement from a finite population
Condition C1 allows convergence rates that are considerably slower than those of parametric estimators [20]. For example, Condition C1 is satisfied when both $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ converge faster than the $M^{1/4}$ rate.
Our proof uses the following von Mises expansion of $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$.
@@
$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$
The first term on the right-hand side of equation (3) is just $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$. The second and third terms will be referred to as, respectively, the ‘empirical process term’ and the ‘remainder term’. We shall prove that these two terms are $o_{p}(M^{-1/2})$ provided that Condition C1 holds. It follows from this that
@@
$\hskip 28.45274pt=\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}+o_{p}(1).$
@@
An important complication when proving that the empirical process and remainder term are $o_{p}(M^{-1/2})$ is that, in general, $\hat{\pi}^{B}_{k}(X)$ is not independent of the data in fold $k$. This is because $\hat{\pi}^{B}_{k}(X)$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{-k}$, $\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{k}$, and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ are, in general, not independent when clusters are sampled without replacement. For example, the number of clusters with $R^{C}_{j}=1$ in fold $k$ must equal $M$ minus the number of clusters with $R^{C}_{j}=1$ in the remaining folds.
4 Simple random sampling of clusters without replacement
Consider the case of simple random sampling of clusters without replacement (so, $\pi^{C}_{1}=\ldots=\pi^{C}_{J}$). As described in Section 2, sampling of individuals within sampled clusters is done using some (possibly multistage) sampling design. Suppose that $M$ and $J$ are both integer multiples of $K$ (the case where $M$ and/or $J$ are not multiples of $K$ is considered as a special case in Section 5).
Randomly partition the $M$ sampled clusters (i.e. those with $R^{C}_{j}=1$) evenly into $K$ folds and, likewise, the JMJ-M unsampled clusters (i.e. those with $R^{C}_{j}=0$) evenly into the $K$ folds, with all such partitions being equally probable. Each fold contains $M/K$ sampled clusters and (JM)/K(J-M)/K unsampled clusters.
This ensures that the conditional distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ is the same as the distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$, and corresponds to simple random sampling of $M/K$ clusters without replacement from the $J/K$ clusters in fold $k$ followed by the original second-stage sampling mechanism for individuals within clusters.
@@ 4 Simple random sampling of clusters without replacement
All this implies that the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ on fold $k$ are conditionally independent of the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$ on the remaining folds given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. Hence, since $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ are calculated using only the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$, the data $\{(R^{A}_{i},R^{B}_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ are conditionally independent of $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. In Appendix A1 we show that this ensures that the empirical process and remainder terms are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 holds.
5 Sampling of clusters without replacement with varying cluster sampling probabilities
Now consider the more general case where there are a small number, $L$, of distinct values of the cluster sampling probability $\pi^{C}_{j}$, which we denote as $\pi^{C(1)},\ldots,\pi^{C(L)}$. Simple random sampling of clusters without replacement is the special case of this with $L=1$. Let $M^{(l)}=M^{(l)}(R^{C}_{1},\ldots,R^{C}_{J})=\sum_{j=1}^{J}R^{C}_{j}\;I(\pi^{C}_{j}=\pi^{C(l)})$ and $J^{(l)}=\sum_{j=1}^{J}I(\pi^{C}_{j}=\pi^{C(l)})$ denote, respectively, the number of sampled clusters and the total number of clusters with cluster sampling probability $\pi^{C(l)}$ ($l=1,\ldots,L$). Clearly, $\sum_{l=1}^{L}J^{(l)}=J$ and $\sum_{l=1}^{L}M^{(l)}=M$. Assume that the cluster sampling mechanism satisfies the following condition:
@@ 5 Sampling of clusters without replacement with varying cluster sampling probabilities
For each of $l=1,\ldots,L$, randomly partition the $M^{(l)}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into $K$ folds, and randomly partition the $J^{(l)}-M^{(l)}$ unsampled clusters evenly into the same $K$ folds. For each $l$, this partition begins by randomly choosing a set of $K\times\lfloor M^{(l)}/K\rfloor$ sampled clusters and a set of $K\times\lfloor(J^{(l)}-M^{(l)})/K\rfloor$ unsampled clusters ($\lfloor x\rfloor$ denotes the integer part of $x$) and evenly partitioning each of these two sets between the folds. This leaves fewer than $K$ sampled clusters and fewer than $K$ unsampled clusters so far unassigned to folds. Each of the unassigned sampled clusters is randomly assigned to the $K$ folds in such a way that no fold receives more than one of these clusters, and the same is done with the unassigned unsampled clusters.
@@
As in Section 4, $\hat{m}_{k}$ is calculated using the $Y$ values of all individuals with $R^{B}=1$ not in fold $k$, and $\hat{\pi}^{B}_{k}$ is calculated using $R^{B}$ values of all the individuals not in fold $k$. However, unlike in Section 4, this calculation of $\hat{\pi}^{B}_{k}$ uses the $R^{A}$ values of only individuals in a subset of the clusters not in fold $k$. We shall refer to this subset as the ‘active’ subset. The reason for using only this subset is to ensure that information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically. We now describe the procedure for choosing the active subset, first for the case where $L=1$ (simple random sampling without replacement) and then for the case where $L>1$.
If $L=1$, then at most $\lceil M/K\rceil$ of the $M$ sampled clusters are in fold $k$ (here, $\lceil x\rceil$ denotes the smallest integer that is greater than or equal to $x$). Hence, at least $M-\lceil M/K\rceil$ sampled clusters must be in the remaining folds. The active subset is chosen by randomly subsampling $M-\lceil M/K\rceil$ of the $M-M_{k}$ sampled clusters not in fold $k$. To compensate for the subsampling, we multiply the $\pi^{A}$ values of individuals not in fold $k$ by $(M-\lceil M/K\rceil)/(M-M/K)$ when calculating $\hat{\pi}^{B}_{k}$.
Now consider the case where $L>1$. Choose a number $0<\delta<1$ and define
@@
($a\wedge b$ denotes the minimum of $a$ and $b$). Choose the active subset by, for each $l=1,\ldots,L$, randomly subsampling $C^{(l)}$ of the $M^{(l)}-M^{(l)}_{k}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ not in fold $k$ To compensate for the subsampling, multiply the $\pi^{A}$ values of individuals with $\pi^{C}=\pi^{C(l)}$ not in fold $k$ by
@@
when calculating $\hat{\pi}^{B}_{k}$. The motivation for calculating $\hat{\pi}^{B}_{k}$ using the $R^{A}$ values only of individuals in the active clusters is that when $M$ is large, $C^{(l)}$ is very likely to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$. If $C^{(l)}$ were guaranteed to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$, then knowing $\hat{\pi}^{B}_{k}$ would tell us nothing about the $(R^{A},R^{B},Y)$ values of individuals in fold $k$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$.
Since, when $M$ is large, $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\approx(M^{(l)}-M^{(l)}_{k})(1-\delta)$, the price we pay to achieve this noninformativeness of $\hat{\pi}^{B}_{k}$ when $M$ is large is the discarding of about $100\delta$% of the individuals with $R^{A}_{i}=1$ when calculating $\hat{\pi}^{B}_{k}$. This motivates us to choose $\delta$ close to zero, e.g. $\delta=0.01$.
In Appendix A2 we prove that when the folds and active subset are chosen in the way described above, the empirical process and remainder terms in the von Mises expansion (equation (3)) are $o_{p}(M^{-1/2})$, as required, provided that Condition C1 is satisfied.

### v2

@@ 3.2 Sampling without replacement from a finite population
$\hat{\theta}_{1}(\underline{\pi}^{B},\underline{m})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}U_{i}(\pi^{B}_{k},m_{k})$
$U_{i}(\pi^{B},m)=\frac{R^{A}_{i}}{\pi^{A}_{i}}m(X_{i})+\frac{R^{B}_{i}}{\pi^{B}(X_{i})}\{Y_{i}-m(X_{i})\}.$
We shall use $\hat{\theta}_{1}(\pi^{B},m)$ as shorthand for $\hat{\theta}_{1}\big((\pi^{B},\ldots,\pi^{B}),(m,\ldots,m)\big)$. Note that $\hat{\theta}_{1}$ is analogous to $\hat{\theta}_{\rm iid}$, but with the weighted mean of $m(X)$ in Sample A substituting for the unknown mean of $m(X)$. We shall assume that $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ satisfy the following condition.
@@
Condition C1 allows convergence rates for $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ considerably slower than those of parametric estimators[21]. For example, it is satisfied when both $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ converge faster than the $M^{1/4}$ rate.
We shall exploit the following von Mises expansion of $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$:
@@
$\hskip 28.45274pt+\sum_{k=1}^{K}\frac{n_{k}}{n}\frac{1}{n_{k}}\sum_{i\in\mathcal{S}_{k}}E\left\{U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})-U_{i}(\pi^{B}_{0},m_{0})\mid\mathcal{F}_{J},\mathcal{S}_{k},\hat{\pi}^{B}_{k},\hat{m}_{k}\right\}.$ (3)
The first term on the right-hand side of equation (3) is just $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$. The second and third terms will be referred to as, respectively, the ‘empirical process term’ and the ‘remainder term’. We shall prove that these two terms are $o_{p}(M^{-1/2})$ provided that Condition C1 holds. It then follows that $\hat{\theta}_{1}(\underline{\hat{\pi}}^{B},\underline{\hat{m}})-\bar{Y}$ has the same asymptotic distribution as $\hat{\theta}_{1}(\pi^{B}_{0},m_{0})-\bar{Y}$, i.e.
@@
$\hskip 28.45274pt=\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{A}_{i}}{\pi^{A}_{i}}m_{0}(X_{i})+\sqrt{M}\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{S}_{k}}\frac{R^{B}_{i}}{\pi^{B}_{0}(X_{i})}\{Y_{i}-m_{0}(X_{i})\}$
$\hskip 42.67912pt-\sqrt{M}\bar{Y}+o_{p}(1).$ (4)
@@
An important complication when proving that the empirical process and remainder term are $o_{p}(M^{-1/2})$ is that, in general, $\hat{\pi}^{B}_{k}$ is not independent of the data in fold $k$. This is because $\hat{\pi}^{B}_{k}(X)$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{-k}$, and $\sum_{i\in\mathcal{S}_{k}}U_{i}(\hat{\pi}^{B}_{k},\hat{m}_{k})$ is calculated using the $R^{A}_{i}$ values of individuals in $\mathcal{S}_{k}$, and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ are not independent when clusters are selected using SWOR. For example, the number of clusters with $R^{C}_{j}=1$ in fold $k$ must equal $M$ minus the number of clusters with $R^{C}_{j}=1$ in the remaining folds. We address this complication in Sections 4 and 5.
4 Simple random sampling without replacement of clusters
First, consider the case of SRSWOR of clusters (so, $\pi^{C}_{1}=\ldots=\pi^{C}_{J}$). As described in Section 2, sampling of individuals within sampled clusters is done using some (possibly multistage) sampling design. Suppose that $M$ and $J$ are both integer multiples of $K$; the case where $M$ and/or $J$ are not multiples of $K$ is considered as a special case in Section 5.
Randomly partition the $M$ sampled clusters (i.e. those with $R^{C}_{j}=1$) evenly into $K$ folds and, likewise, the JMJ-M unsampled clusters (i.e. those with $R^{C}_{j}=0$) evenly into the $K$ folds, with all such partitions being equally probable. Thus, each fold contains $M/K$ sampled clusters and (JM)/K(J-M)/K unsampled clusters.
This ensures that the conditional distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$, $\mathcal{S}_{k}$ and $\{R^{A}_{i}:\;i\in\mathcal{S}_{-k}\}$ is the same as the distribution of $\{R^{A}_{i}:\;i\in\mathcal{S}_{k}\}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$, and corresponds to SRSWOR of $M/K$ clusters from the $J/K$ clusters in fold $k$ followed by the original second-stage sampling mechanism for individuals within clusters.
@@ 4 Simple random sampling without replacement of clusters
All this implies that the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ on fold $k$ are conditionally independent of the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$ on the remaining folds given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. Hence, since $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ are calculated using only the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{-k}\}$, the data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\in\mathcal{S}_{k}\}$ are conditionally independent of $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. In Appendix A1 we show that this ensures that the empirical process and remainder terms are $o_{p}(M^{-1/2})$, provided that Condition C1 holds.
5 Varying cluster sampling probabilities
Now consider the more general case where there are $L$ distinct values of the cluster sampling probability $\pi^{C}_{j}$, which we denote as $\pi^{C(1)},\ldots,\pi^{C(L)}$. Assume that $L$ is small relative to $M$. SRSWOR of clusters is the special case of this with $L=1$. Let $M^{(l)}=M^{(l)}(R^{C}_{1},\ldots,R^{C}_{J})=\sum_{j=1}^{J}R^{C}_{j}\;I(\pi^{C}_{j}=\pi^{C(l)})$ and $J^{(l)}=\sum_{j=1}^{J}I(\pi^{C}_{j}=\pi^{C(l)})$ denote, respectively, the number of sampled clusters and the total number of clusters with cluster sampling probability $\pi^{C(l)}$ ($l=1,\ldots,L$). Clearly, $\sum_{l=1}^{L}J^{(l)}=J$ and $\sum_{l=1}^{L}M^{(l)}=M$. Assume that the cluster sampling mechanism satisfies the following symmetry condition:
@@ 5 Varying cluster sampling probabilities
For each of $l=1,\ldots,L$, randomly partition the $M^{(l)}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into $K$ folds, and randomly partition the $J^{(l)}-M^{(l)}$ unsampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ evenly into the same $K$ folds. For each $l$, this partition begins by randomly choosing a set of $K\times\lfloor M^{(l)}/K\rfloor$ sampled clusters ($\lfloor x\rfloor$ denotes the integer part of $x$) and evenly partitioning these between the folds. This leaves fewer than $K$ sampled clusters so far unassigned to folds. Each of these is randomly assigned to the $K$ folds in such a way that no fold receives more than one of them. The same is done with the unassigned unsampled clusters.
@@
As in Section 4, $\hat{m}_{k}$ is calculated using the $Y$ and $X$ values of all individuals with $R^{B}=1$ not in fold $k$, and $\hat{\pi}^{B}_{k}$ uses the data $\{(R^{B}_{i},R^{B}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ on all individuals not in fold $k$. However, unlike in Section 4, where $\hat{\pi}^{B}_{k}$ used the data $\{(R^{A}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ on all individuals not in fold $k$, here we discard some of these data when calculating $\hat{\pi}^{B}_{k}$. Specifically, we only use data $(R^{A},R^{A}\pi^{A},R^{A}X)$ on those individuals in Sample A who belong to a subset of the clusters not in fold $k$. We refer to this subset of clusters as the ‘active’ subset. In effect, we temporarily set $R^{A}=R^{A}X=0$ for any individual in Sample A and not in fold $k$ who does not belong to an active cluster while calculating $\hat{\pi}^{B}_{k}$. As explained below, the reason for only using this subset is to ensure that information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically.
To choose the active subset, first choose a number $0<\delta<1$ and define
@@
(here, $a\wedge b$ denotes the minimum of $a$ and $b$). For each $l=1,\ldots,L$, randomly subsample $C^{(l)}$ of the $M^{(l)}-M^{(l)}_{k}$ sampled clusters with $\pi^{C}_{j}=\pi^{C(l)}$ not in fold $k$. The set of $\sum_{l=1}^{L}M^{(l)}-M^{(l)}_{k}$ subsampled clusters is the active subset. To compensate for the subsampling, we temporarily multiply the $\pi^{A}$ values of individuals with $\pi^{C}=\pi^{C(l)}$ not in fold $k$ by
@@
when calculating $\hat{\pi}^{B}_{k}$.
The motivation for calculating $\hat{\pi}^{B}_{k}$ using the $(R^{A},R^{A}X)$ values only of individuals in the active clusters is that when $M$ is large, $C^{(l)}$ is very likely to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$. If $C^{(l)}$ were guaranteed to equal $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor$ for all $l$, then knowing $\hat{\pi}^{B}_{k}$ would tell us nothing about the $(R^{A},R^{B},Y)$ values of individuals in fold $k$ given $\mathcal{F}_{J}$ and $\mathcal{S}_{k}$. Hence, information about the $R^{A}$ values of individuals in fold $k$ provided by the value of $\hat{\pi}^{B}_{k}$ vanishes asymptotically.
When $M$ is large, $\lfloor\pi^{C(l)}(1-\delta)(J^{(l)}-J^{(l)}_{k})\rfloor\approx(M^{(l)}-M^{(l)}_{k})(1-\delta)$. So, the price we pay to achieve this noninformativeness of $\hat{\pi}^{B}_{k}$ when $M$ is large is the discarding of about $100\delta$% of the individuals with $R^{A}_{i}=1$ when calculating $\hat{\pi}^{B}_{k}$. This motivates us to choose $\delta$ close to zero, e.g. $\delta=0.01$.
In Appendix A2 we prove that when the folds and active subset are chosen in the way described above, the empirical process and remainder terms in the von Mises expansion (equation (3)) are $o_{p}(M^{-1/2})$, provided that Condition C1 is satisfied.
In the special case of $L=1$ (SRSWOR), we can instead choose the active subset to be a random subset of $M-\lceil M/K\rceil$ of the $M-M_{k}$ sampled clusters not in fold $k$ ($\lceil x\rceil$ denotes the smallest integer greater than or equal to $x$). This ensures noninformativeness of $\hat{\pi}^{B}_{k}$, because there must be at least $M-\lceil M/K\rceil$ sampled clusters not in fold $k$, and ensures that at most K1K-1 sampled clusters are excluded from the active subset. To compensate for the subsampling, we temporarily multiply the $\pi^{A}$ values of individuals not in fold $k$ by $(M-\lceil M/K\rceil)/(M-M/K)$ when calculating $\hat{\pi}^{B}_{k}$. Clearly, if $M$ is an integer multiple of $K$, then all sampled clusters not in fold $k$ are in the active subset, i.e. there is no subsampling.